# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 4.5.1 Solution to the Regularized Dyson Equation
For every $\tau\in\mathbb{R}_{>0}$, the existence of a unique $M\in\mathcal{M}_{+}$ satisfying (5) for every $z\in\mathbb{H}$ follows directly from [37]. At a high level, the proof of the existence of a solution to the RDEL (5) proceeds by demonstrating that the map $M\mapsto\mathcal{F}^{(\tau)}(M)$ is strictly holomorphic on bounded domains of $\mathcal{M}_{+}$. By the Earle-Hamilton fixed-point theorem, strict holomorphicity implies that the mapping is contractive with respect to the CRF-pseudometric. For further details on the CRF-pseudometric, we refer the reader to Section A.1.
@@
for every $b>0$. Indeed, for every $b\in\mathbb{R}_{>0}$, $\mathcal{M}_{b}$ is a domain in the Banach space of matrix-valued bounded holomorphic functions on $\mathbb{H}$ with the canonical supremum norm. Also, $\mathcal{A}_{b}$ is a domain in the Banach space of complex symmetric $\ell\times\ell$ matrices with the operator norm. Using the result above, we can show that $\mathcal{F}^{(\tau)}$ is indeed a strict holomorphic function on $\mathcal{A}_{b}$ for every $\tau\in\mathbb{R}_{>0}$. The following lemma is a direct adaptation of [37, Proposition 3.2].
@@
The existence of a unique solution to (5) then follows directly from an application of the Earle-Hamilton fixed-point theorem [23]. Indeed, for every $b\in\mathbb{R}_{>0}$ large enough, $\mathcal{F}^{(\tau)}$ has exactly one fixed point on $\mathcal{M}_{b}$. Since $\mathcal{M}_{+}=\bigcup_{b\in\mathbb{R}_{>0}}\mathcal{M}_{b}$, we obtain the following result.
Proposition 4.15 ([37, Theorem 2.1]).
@@
In what follows, we will denote the unique solution of the RDEL with $\tau\in\mathbb{R}_{>0}$ by $M^{(\tau)}$. While not explicitly stated, the analyticity of $M^{(\tau)}(z)$ in $\tau$ can be inferred using an implicit function theorem, as demonstrated in [26, Theorem 2.14]. We state this result in the following lemma but omit the proof, directing the reader to the aforementioned reference for further details.
@@ 4.5.2 Solution to the Dyson Equation for Linearizations
[…]
@@
for some analytic function $f$. By induction hypothesis, $f(M_{0},M_{1},\ldots,M_{k})$ may be expressed as an analytic function of $\{C_{j}\,:\,j\in\mathbb{N}_{\geq 0}\}$. Furthermore, since $C_{0}=M_{\infty}$ is 00 everywhere outside its lower $d\times d$ block,
@@ 5.2 Stability
Before stating the first lemma, we recall from (18) and (4.14) that $\mathcal{A}_{b}:=\mathcal{A}_{+}\cap\{W\in\mathbb{C}^{\ell\times\ell}\,:\,\|W\|<b\}$ is a domain in the Banach space of $\ell\times\ell$ complex matrices for every $b\in\mathbb{R}_{>0}$. The CRF-pseudometric is a crucial tool because $\mathcal{F}^{(\tau)}$ is a strict contraction on $\mathcal{A}_{b}$ with respect to the CRF-pseudometric. This can be observed and quantified by combining Lemma 4.14 with the proof of the Earle-Hamilton fixed-point theorem [23].
@@
Let $X\in\mathcal{A}_{b}$ and define $\mathcal{G}:Y\in\mathcal{A}_{b}\mapsto\mathcal{F}^{(\tau)}(Y)+\delta(\mathcal{F}^{(\tau)}(Y)-\mathcal{F}^{(\tau)}(X))$. By Lemma 4.14, $\|\mathcal{F}^{(\tau)}(Y)\|\leq\tau^{-1}$ and $\Im[\mathcal{F}^{(\tau)}(Y)]\succ\tau m_{b}^{-2}$ for every $Y\in\mathcal{A}_{b}$. Hence, $\|\mathcal{G}(Y)\|\leq\tau^{-1}+2\delta\tau^{-1}$ and $\Im[\mathcal{G}(Y)]\succeq(1+\delta)\Im[\mathcal{F}^{(\tau)}(Y)]-\delta\|\mathcal{F}^{(\tau)}(X)\|\succ(1+\delta)\tau m_{b}^{-2}-\delta\tau^{-1}$. By our choice of $\delta$ and $b$, $\|\mathcal{G}(Y)\|<b$ and $\Im[G(Y)]\succ 0$ for every $Y\in\mathcal{A}_{b}$. In fact, $\mathcal{G}$ is a strict holomorphic function on $\mathcal{A}_{b}$. By the proof of the Earle-Hamilton fixed-point theorem in [35, Theorem 3.1] and [34, Theorem 4], we obtain $\rho(\mathcal{F}^{(\tau)}(X),\mathcal{F}^{(\tau)}(Y))\leq(1+\delta)^{-1}\rho(X,Y)$. ∎
@@
Let $m_{b}$ be defined as in the proof of Lemma 5.3 and recall that $\|M^{(\tau)}\|\leq\tau^{-1}$ as well as $\Im[M^{(\tau)}]\succ\tau m_{b}^{-2}$. In particular, $M^{(\tau)},F\in\mathcal{A}_{b}$. Define the holomorphic function $f:X\in\mathcal{A}_{b}\mapsto\mathrm{tr}(U(X-M^{(\tau)}))(b+\tau^{-1})^{-1}\in\mathbb{D}$. By Lemma A.9, $|\mathrm{tr}(U(X-M^{(\tau)}))|\leq\|X-M^{(\tau)}\|<b+\tau^{-1}$, which ensures that $f$ is well-defined. By Propositions A.1 and 26,
@@
By Corollary 5.5, $\|M^{(\tau)}-F^{(\tau)}\|\leq(b+\tau^{-1})\tanh(\rho(M^{(\tau)},F^{(\tau)}))$. Recursively define a sequence $\{M_{k}\,:\,k\in\mathbb{N}_{0}\}\subseteq\mathcal{A}_{b}$ such that $M_{0}=F^{(\tau)}$ and $M_{k}=\mathcal{F}^{(\tau)}(M_{k-1})$ for every $k\in\mathbb{N}$. By lemmas 5.3 and 5.6,
@@
To summarize, if $F^{(\tau)}\in\mathcal{A}_{\tau^{-1}+2\tau}$ approximately solves (5) up to an additive perturbation term $D^{(\tau)}$ in the sense of (19), and $\|\mathcal{F}^{(\tau)}(F^{(\tau)})D^{(\tau)}\|\leq\tau^{-1}\|D^{(\tau)}\|$ vanishes as $\ell\to\infty$, then $\|M^{(\tau)}-F^{(\tau)}\|$ converges to 00 as $\ell\to\infty$. Here, $M^{(\tau)}$ refers to the unique solution to (5). Notably, if we take $F^{(\tau)}=\mathbb{E}(L-z\Lambda-i\tau I_{\ell})^{-1}$, then we obtain a way to establish pointwise convergence in $z$ and $\tau$. Utilizing conditions (C1) and (C2) as well as lemmas 5.1 and 5.2, along with a diagonalization argument, we can establish the following result.
@@
The proof of Corollary 5.8 highlights a contrast between $\tau$ and $\ell$ regarding their impact on convergence behavior. On one hand, as $\tau$ approaches 00, the solution to the regularized Dyson equation (5) converges to the solution of the Dyson equation (2). However, the solution to the Dyson equation possesses less desirable properties compared to the regularized solution, notably because the imaginary part is not guaranteed to be positive definite. On the other hand, as $\ell$ approaches $\infty$, we establish stability for fixed $\tau$ by controlling the magnitude of the error $\epsilon_{\tau}$. We leverage this stability to establish convergence between the expected regularized pseudo-resolvent and the solution to the DEL. Therefore, we need to allow $\tau$ to approach 00 slowly enough as $\ell\to\infty$ to preserve the stability property.
@@ 5.3 Perturbation
By the Gaussian concentration inequality for Lipschitz functions [58, Theorem 2.1.12] and [44, Proposition 1.10], there exists some absolute constant $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that
@@
for all unit vectors $u,v\in\mathbb{C}^{\ell}$. Suppose that $\epsilon\in(0,2^{-3/2})$ and let $\mathcal{N}$ be an $\epsilon$-net for the unit ball of $\ell$-dimensional real vectors. Then, given $u\in\mathbb{C}^{\ell}$, we may find $v_{1},v_{2}\in\mathcal{N}$ such that $\|u-v_{1}-iv_{2}\|^{2}=\|\Re[u]-v_{1}\|^{2}+\|\Im[u]-v_{2}\|^{2}\leq 2\epsilon^{2}$. In particular, $\mathcal{N}+i\mathcal{N}:=\{v_{1}+iv_{2}\,:\,v_{1},v_{2}\in\mathcal{N}\}$ forms a $\sqrt{2}\epsilon$-net for the unit sphere of $\ell$-dimensional complex unitary vectors. By [60, Corollary 4.2.13], $|\mathcal{N}+i\mathcal{N}|\leq(2\epsilon^{-1}+1)^{2\ell}$.
@@
As a direct outcome of lemmas 5.10 and 4.3, it follows that as the dimension $\ell$ tends towards infinity, $\|D^{(\tau)}\|$ diminishes, provided that $\lim_{\ell\to\infty}\sqrt{\ell}\lambda=0$, $\lim_{\ell\to\infty}\|\mathcal{S}\|=0$, and $\lim_{\ell\to\infty}\|\Delta(L,\tau)\|=0$ for every $\tau\in\mathbb{R}_{>0}$ sufficiently small. Before we proceed, we highlight two significant observations.
@@ 5.4 Concentration
Let $g_{1},g_{2}\in\mathbb{R}^{\gamma}$. Then, by Cauchy-Schwarz’s inequality $|\mathrm{tr}(U((L(g_{1})-z\Lambda)^{-1}-(L(g_{2})-z\Lambda)^{-1}))|\leq\|U\|_{F}\|(L(g_{1})-z\Lambda)^{-1}-(L(g_{2})-z\Lambda)^{-1}\|_{F}\leq\lambda\|g_{1}-g_{2}\|$. By Gaussian concentration inequality for Lipschitz functions [58, 44], there exists absolute constants $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that $\mathbb{P}(|\mathrm{tr}(U((L-z\Lambda)^{-1}-\mathbb{E}(L-z\Lambda)^{-1}))|\geq x)\leq c_{1}e^{-c_{2}x^{2}/\lambda^{2}}\leq c_{1}e^{-c_{3}x^{2}\ell^{2r}}$ for every $x\in\mathbb{R}_{>0}$ and $\ell\in\mathbb{N}$. Here, $c_{3}\in\mathbb{R}_{>0}$ is some constant satisfying $0<c_{3}\leq c_{2}/(\lambda\ell^{r})^{2}$ for every $\ell\in\mathbb{N}$ large enough. In particular, for any $\epsilon\in\mathbb{R}_{>0}$, $\sum_{\ell=1}^{\infty}\mathbb{P}(|\mathrm{tr}(U((L-z\Lambda)^{-1}-\mathbb{E}(L-z\Lambda)^{-1}))|\geq\epsilon)\leq c_{1}\sum_{\ell=1}^{\infty}e^{-c_{3}\epsilon^{2}\ell^{2r}}$. By the integral test, the series $\sum_{\ell=1}^{\infty}e^{-c_{3}\epsilon^{2}\ell^{2r}}$ converges, and the result follows from the Borel-Cantelli lemma. ∎
@@ 6 Proof of Asymptotic Equivalence for Random Features
$E_{\text{test}}=\|\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}y\|^{2}-2\hat{y}^{T}\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}y+\|\hat{y}\|^{2}.$
Each term in the above equations is a bilinear form, aligning well with the framework of deterministic equivalence. In order to apply our framework, we have to find a linearization such that a matrix of interest is contained in one of the block of the inverse. To this end, let $\ell=n_{\text{train}}+d+2n_{\text{test}}$ and consider the linearization
$L=\begin{bmatrix}\delta I_{n_{\text{train}}}&A&0_{n_{\text{train}}\times n_{\text{test}}}&0_{n_{\text{train}}\times n_{\text{test}}}\\ A^{T}&-I_{d\times d}&0_{d\times n_{\text{test}}}&\hat{A}^{T}\\ 0_{n_{\text{test}}\times n_{\text{train}}}&0_{n_{\text{test}}\times d}&0_{n_{\text{test}}\times n_{\text{test}}}&-I_{n_{\text{test}}}\\ 0_{n_{\text{test}}\times n_{\text{train}}}&\hat{A}&-I_{n_{\text{test}}}&0_{n_{\text{test}}\times n_{\text{test}}}\end{bmatrix}\in\mathbb{R}^{\ell\times\ell}.$ (22)
@@
$(L-z\Lambda)^{-1}=\begin{bmatrix}R&(1+z)^{-1}RA&(1+z)^{-1}RA\hat{A}^{T}&0\\ (1+z)^{-1}A^{T}R&\bar{R}&\bar{R}\hat{A}^{T}&0\\ (1+z)^{-1}\hat{A}A^{T}R&\hat{A}\bar{R}&\hat{A}\bar{R}\hat{A}^{T}&-I_{n_{\text{test}}}\\ 0&0&-I_{n_{\text{test}}}&0\end{bmatrix}.$
Here $R:=((1+z)^{-1}AA^{T}+(\delta-z)I_{n_{\text{train}}})^{-1}$ represents a resolvent and $\bar{R}:=-((1+z)I_{d}+(\delta-z)^{-1}A^{T}A)^{-1}$ is a co-resolvent. Indeed, $\lim_{z\to 0}(L-z\Lambda)^{-1}_{3,1}=\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$ holds one of the relevant expression for which we want to find a deterministic equivalent. Therefore, it suffices to find a deterministic equivalent for the pseudo-resolvent $(L-z\Lambda)^{-1}$ and take the spectral parameter to zero in order to recover a deterministic equivalent for $\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$.

### v4

@@ 4.5.1 Solution to the Regularized Dyson Equation
For every $\tau\in\mathbb{R}_{>0}$, the existence of a unique $M\in\mathcal{M}_{+}$ satisfying (5) for every $z\in\mathbb{H}$ follows directly from [helton2007operatorvalued]. At a high level, the proof of the existence of a solution to the RDEL (5) proceeds by demonstrating that the map $M\mapsto\mathcal{F}^{(\tau)}(M)$ is strictly holomorphic on bounded domains of $\mathcal{M}_{+}$. By the Earle-Hamilton fixed-point theorem, strict holomorphicity implies that the mapping is contractive with respect to the CRF-pseudometric. For further details on the CRF-pseudometric, we refer the reader to Section A.1.
@@
for every $b>0$. Indeed, for every $\in\mathbb{R}_{>0}, b\mathcal{M}_{b}$ is a domain in the Banach space of matrix-valued bounded holomorphic functions on $\mathbb{H}$ with the canonical supremum norm. Also, $\mathcal{A}_{b}$ is a domain in the Banach space of complex symmetric $\ell\times\ell$ matrices with the operator norm. Using the result above, we can show that $\mathcal{F}^{(\tau)}$ is indeed a strict holomorphic function on $\mathcal{A}_{b}$ for every $\tau\in\mathbb{R}_{>0}$. The following lemma is a direct adaptation of [helton2007operatorvalued, Proposition 3.2].
@@
The existence of a unique solution to (5) then follows directly from an application of the Earle-Hamilton fixed-point theorem [earl_hamilton]. Indeed, for every $b\in\mathbb{R}_{>0}$ large enough, $\mathcal{F}^{(\tau)}$ has exactly one fixed point on $\mathcal{M}_{b}$. Since $\mathcal{M}_{+}=\bigcup_{b\in\mathbb{R}_{>0}}\mathcal{M}_{b}$, we obtain the following result.
Proposition 4.1 ([helton2007operatorvalued, Theorem 2.1]).
@@
In what follows, we will denote the unique solution of the RDEL with $\tau\in\mathbb{R}_{>0}$ by $M^{(\tau)}$. While not explicitly stated, the analyticity of $M^{(\tau)}(z)$ in $\tau$ can be inferred using an implicit function theorem, as demonstrated in [nemish_local_2020, Theorem 2.14]. We state this result in the following lemma but omit the proof, directing the reader to the aforementioned reference for further details.
@@ 4.5.2 Solution to the Dyson Equation for Linearizations
We now proceed inductively. Assume that there exists $m\in\mathbb{N}$ and a collection of compactly supported measures $\{\omega_{j}\,:\,1\leq j\leq m\}$ such that $v_{j}^{\ast}\Omega_{1,1}^{(\tau_{m,k})}v_{j}$ converges weakly to $\omega_{j}$ for all $1\leq j\leq m$ as $k$ approaches infinity. By Corollary 4.3 and Prokhorov’s theorem, there exists a probability measure $\omega_{m+1}$ and a subsequence $\{\tau_{m+1,k}\,:\,k\in\mathbb{N}\}\subseteq\{\tau_{m,k}\,:\,k\in\mathbb{N}\}$ such that $v_{m+1}^{\ast}\Omega_{1,1}^{(\tau_{m+1,k})}v_{m+1}$ converges weakly to $\omega_{m+1}$ as $k$ approaches infinity. Also, by construction of the subsequence, $v_{j}^{\ast}\Omega_{1,1}^{(\tau_{m+1,k})}v_{j}$ converges weakly to $\omega_{j}$ for all $1\leq j\leq m+1$ as $k$ approaches infinity.
@@
for some analytic function $f$. By induction hypothesis, $f(M_{0},M_{1},\ldots,M_{k})$ may be expressed as an analytic function of $\{C_{j}\,:\,j\in\mathbb{N}_{\geq 0}\}$. Furthermore, since $C_{0}=M_{\infty}$ is 0 everywhere outside its lower $d\times d$ block,
@@ 5.2 Stability
Before stating the first lemma, we recall from (18) and (4.10) that $\mathcal{A}_{b}:=\mathcal{A}_{+}\cap\{W\in\mathbb{C}^{\ell\times\ell}\,:\,\|W\|<b\}$ is a domain in the Banach space of $\ell\times\ell$ complex matrices for every $b\in\mathbb{R}_{>0}$. The CRF-pseudometric is a crucial tool because $\mathcal{F}^{(\tau)}$ is a strict contraction on $\mathcal{A}_{b}$ with respect to the CRF-pseudometric. This can be observed and quantified by combining Lemma 4.10 with the proof of the Earle-Hamilton fixed-point theorem [earl_hamilton].
@@
Let $X\in\mathcal{A}_{b}$ and define $\mathcal{G}:Y\in\mathcal{A}_{b}\mapsto\mathcal{F}^{(\tau)}(Y)+\delta(\mathcal{F}^{(\tau)}(Y)-\mathcal{F}^{(\tau)}(X))$. By Lemma 4.10, $\|\mathcal{F}^{(\tau)}(Y)\|\leq\tau^{-1}$ and $\Im[\mathcal{F}^{(\tau)}(Y)]\succ\tau m_{b}^{-2}$ for every $Y\in\mathcal{A}_{b}$. Hence, $\|\mathcal{G}(Y)\|\leq\tau^{-1}+2\delta\tau^{-1}$ and $\Im[\mathcal{G}(Y)]\succeq(1+\delta)\Im[\mathcal{F}^{(\tau)}(Y)]-\delta\|\mathcal{F}^{(\tau)}(X)\|\succ(1+\delta)\tau m_{b}^{-2}-\delta\tau^{-1}$. By our choice of $\delta$ and $b$, $\|\mathcal{G}(Y)\|<b$ and $\Im[G(Y)]\succ 0$ for every $Y\in\mathcal{A}_{b}$. In fact, $\mathcal{G}$ is a strict holomorphic function on $\mathcal{A}_{b}$. By the proof of the Earle-Hamilton fixed-point theorem in [harris_fixed_2003, Theorem 3.1] and [HARRIS1979345, Theorem 4], we obtain $\rho(\mathcal{F}^{(\tau)}(X),\mathcal{F}^{(\tau)}(Y))\leq(1+\delta)^{-1}\rho(X,Y)$. ∎
@@
Let $m_{b}$ be defined as in the proof of Lemma 5.3 and recall that $\|M^{(\tau)}\|\leq\tau^{-1}$ as well as $\Im[M^{(\tau)}]\succ\tau m_{b}^{-2}$. In particular, $M^{(\tau)},F\in\mathcal{A}_{b}$. Define the holomorphic function $f:X\in\mathcal{A}_{b}\mapsto\mathrm{tr}(U(X-M^{(\tau)}))(b+\tau^{-1})^{-1}\in\mathbb{D}$. By Lemma A.8, $|\mathrm{tr}(U(X-M^{(\tau)}))|\leq\|X-M^{(\tau)}\|<b+\tau^{-1}$, which ensures that $f$ is well-defined. By Proposition A.1 and (26),
@@
By Corollary 5.1, $\|M^{(\tau)}-F^{(\tau)}\|\leq(b+\tau^{-1})\tanh(\rho(M^{(\tau)},F^{(\tau)}))$. Recursively define a sequence $\{M_{k}\,:\,k\in\mathbb{N}_{0}\}\subseteq\mathcal{A}_{b}$ such that $M_{0}=F^{(\tau)}$ and $M_{k}=\mathcal{F}^{(\tau)}(M_{k-1})$ for every $k\in\mathbb{N}$. By lemmas 5.3 and 5.5,
@@
To summarize, if $F^{(\tau)}\in\mathcal{A}_{\tau^{-1}+2\tau}$ approximately solves (5) up to an additive perturbation term $D^{(\tau)}$ in the sense of (19), and $\|\mathcal{F}^{(\tau)}(F^{(\tau)})D^{(\tau)}\|\leq\tau^{-1}\|D^{(\tau)}\|$ vanishes as $\ell\to\infty$, then $\|M^{(\tau)}-F^{(\tau)}\|$ converges to 0 as $\ell\to\infty$. Here, $M^{(\tau)}$ refers to the unique solution to (5). Notably, if we take $F^{(\tau)}=\mathbb{E}(L-z\Lambda-i\tau I_{\ell})^{-1}$, then we obtain a way to establish pointwise convergence in $z$ and $\tau$. Utilizing conditions (C1) and (C2) as well as lemmas 5.1 and 5.2, along with a diagonalization argument, we can establish the following result.
@@
The proof of Corollary 5.2 highlights a contrast between $\tau$ and $\ell$ regarding their impact on convergence behavior. On one hand, as $\tau$ approaches 0, the solution to the regularized Dyson equation (5) converges to the solution of the Dyson equation (2). However, the solution to the Dyson equation possesses less desirable properties compared to the regularized solution, notably because the imaginary part is not guaranteed to be positive definite. On the other hand, as $\ell$ approaches $\infty$, we establish stability for fixed $\tau$ by controlling the magnitude of the error $\epsilon_{\tau}$. We leverage this stability to establish convergence between the expected regularized pseudo-resolvent and the solution to the DEL. Therefore, we need to allow $\tau$ to approach 0 slowly enough as $\ell\to\infty$ to preserve the stability property.
@@ 5.3 Perturbation
By the Gaussian concentration inequality for Lipschitz functions [tao_rmt, Theorem 2.1.12] and [ledoux_concentration_2001, Proposition 1.10], there exists some absolute constant $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that
@@
for all unit vectors $u,v\in\mathbb{C}^{\ell}$. Suppose that $\epsilon\in(0,2^{-3/2})$ and let $\mathcal{N}$ be an $\epsilon$-net for the unit ball of $\ell$-dimensional real vectors. Then, given $u\in\mathbb{C}^{\ell}$, we may find $v_{1},v_{2}\in\mathcal{N}$ such that $\|u-v_{1}-iv_{2}\|^{2}=\|\Re[u]-v_{1}\|^{2}+\|\Im[u]-v_{2}\|^{2}\leq 2\epsilon^{2}$. In particular, $\mathcal{N}+i\mathcal{N}:=\{v_{1}+iv_{2}\,:\,v_{1},v_{2}\in\mathcal{N}\}$ forms a $\sqrt{2}\epsilon$-net for the unit sphere of $\ell$-dimensional complex unitary vectors. By [vershynin_2018, Corollary 4.2.13], $|\mathcal{N}+i\mathcal{N}|\leq(2\epsilon^{-1}+1)^{2\ell}$.
@@
As a direct outcome of lemmas 5.8 and 4.2, it follows that as the dimension $\ell$ tends towards infinity, $\|D^{(\tau)}\|$ diminishes, provided that $\lim_{\ell\to\infty}\sqrt{\ell}\lambda=0$, $\lim_{\ell\to\infty}\|\mathcal{S}\|=0$, and $\lim_{\ell\to\infty}\|\Delta(L,\tau)\|=0$ for every $\tau\in\mathbb{R}_{>0}$ sufficiently small. Before we proceed, we highlight two significant observations.
@@ 5.4 Concentration
Let $g_{1},g_{2}\in\mathbb{R}^{\gamma}$. Then, by Cauchy-Schwarz’s inequality $|\mathrm{tr}(U((L(g_{1})-z\Lambda)^{-1}-(L(g_{2})-z\Lambda)^{-1}))|\leq\|U\|_{F}\|(L(g_{1})-z\Lambda)^{-1}-(L(g_{2})-z\Lambda)^{-1}\|_{F}\leq\lambda\|g_{1}-g_{2}\|$. By Gaussian concentration inequality for Lipschitz functions [tao_rmt, ledoux_concentration_2001], there exists absolute constants $c_{1},c_{2}\in\mathbb{R}_{>0}$ such that $\mathbb{P}(|\mathrm{tr}(U((L-z\Lambda)^{-1}-\mathbb{E}(L-z\Lambda)^{-1}))|\geq x)\leq c_{1}e^{-c_{2}x^{2}/\lambda^{2}}\leq c_{1}e^{-c_{3}x^{2}\ell^{2r}}$ for every $x\in\mathbb{R}_{>0}$ and $\ell\in\mathbb{N}$. Here, $c_{3}\in\mathbb{R}_{>0}$ is some constant satisfying $0<c_{3}\leq c_{2}/(\lambda\ell^{r})^{2}$ for every $\ell\in\mathbb{N}$ large enough. In particular, for any $\epsilon\in\mathbb{R}_{>0}$, $\sum_{\ell=1}^{\infty}\mathbb{P}(|\mathrm{tr}(U((L-z\Lambda)^{-1}-\mathbb{E}(L-z\Lambda)^{-1}))|\geq\epsilon)\leq c_{1}\sum_{\ell=1}^{\infty}e^{-c_{3}\epsilon^{2}\ell^{2r}}$. By the integral test, the series $\sum_{\ell=1}^{\infty}e^{-c_{3}\epsilon^{2}\ell^{2r}}$ converges, and the result follows from the Borel-Cantelli lemma. ∎
@@ 6 Proof of Asymptotic Equivalence for Random Features
$E_{\text{test}}=\|\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}y\|^{2}-2\hat{y}^{\top}\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}y+\|\hat{y}\|^{2}.$
Each term in the above equations is a bilinear form, aligning well with the framework of deterministic equivalence. In order to apply our framework, we have to find a linearization such that a matrix of interest is contained in one of the blocks of the inverse. To this end, let $\ell=n_{\text{train}}+d+2n_{\text{test}}$ and consider the linearization
$L=\begin{bmatrix}\delta I_{n_{\text{train}}}&A&0_{n_{\text{train}}\times n_{\text{test}}}&0_{n_{\text{train}}\times n_{\text{test}}}\\ A^{\top}&-I_{d\times d}&0_{d\times n_{\text{test}}}&\hat{A}^{\top}\\ 0_{n_{\text{test}}\times n_{\text{train}}}&0_{n_{\text{test}}\times d}&0_{n_{\text{test}}\times n_{\text{test}}}&-I_{n_{\text{test}}}\\ 0_{n_{\text{test}}\times n_{\text{train}}}&\hat{A}&-I_{n_{\text{test}}}&0_{n_{\text{test}}\times n_{\text{test}}}\end{bmatrix}\in\mathbb{R}^{\ell\times\ell}$. (22)
@@
$(L-z\Lambda)^{-1}=\begin{bmatrix}R&(1+z)^{-1}RA&(1+z)^{-1}RA\hat{A}^{\top}&0\\ (1+z)^{-1}A^{\top}R&\bar{R}&\bar{R}\hat{A}^{\top}&0\\ (1+z)^{-1}\hat{A}A^{\top}R&\hat{A}\bar{R}&\hat{A}\bar{R}\hat{A}^{\top}&-I_{n_{\text{test}}}\\ 0&0&-I_{n_{\text{test}}}&0\end{bmatrix}$.
Here $R:=((1+z)^{-1}AA^{\top}+(\delta-z)I_{n_{\text{train}}})^{-1}$ represents a resolvent and $\bar{R}:=-((1+z)I_{d}+(\delta-z)^{-1}A^{\top}A)^{-1}$ is a co-resolvent. Indeed, $\lim_{z\to 0}(L-z\Lambda)^{-1}_{3,1}=\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$ holds one of the relevant expression for which we want to find a deterministic equivalent. Therefore, it suffices to find a deterministic equivalent for the pseudo-resolvent $(L-z\Lambda)^{-1}$ and take the spectral parameter to zero in order to recover a deterministic equivalent for $\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$.